OpenStudy (anonymous):

f you rent a car for one day and drive it for 100 miles, the cost is $40.00. If you drive it 220 miles, the cost is $46.00. Choose the linear equation that would represent the data points.

OpenStudy (anonymous):

(100,40),(220,46) are the two points. find the equation of the line between them

OpenStudy (anonymous):

slope is \[\frac{46-40}{220-100}=\frac{6}{120}=.05\] so you are charged 5 cents a mile

OpenStudy (anonymous):

flat fee is the y intercept. you can use \[y-y_1=m(x-x_1)\] or \[y-40=.05(x-100)\] \[y-40=.05x-5\] \[y=.05x+35\] meaning you are charged $35 to rent the car and 5 cents a mile